Grade 11General Physics

If there is a positive error of 50% in the measurement of velocity of a body then the error in the measurement of kinetic energy is? 100% or 125%

Prachi
10 Years agoGrade 11

2 Answers

Vikas TU
10 Years ago
K.E. = 0.5mv^2
dk = 0.5m*2v*dv
 
k.e is  propr. to square of v
error in K.E = 0.5^2v = 0.25K
therfore,
= k + 0.25k
= 1.25k
= 125%
 
Ashwin
9 Years ago
KE =1/2mv²∆KE/KE×100=2∆v/v×100since ∆v/v×100=50%by substituting the value above we get∆KE/KE×100=100%